0.8 divided by 1/4=(the answers are.....16/5, 3.08, 0.2, or none of these)
(please, please, please, respond quickly thank you!!!!!!!!!!!!!!!!!!!!!!)

Found 3 solutions by DWL, BadLandsBarn, Alan3354:
Answer by DWL(56)

You must convert the .8 decimal to fraction. .8 = 4/5
then set up the equation 4/5 divided by 1/4
Dividing by a fraction is the same as multiplying by it's reciprocal:
4/5 times 4/1
4 times 4 = 16
5 times 1 = 5
answer = 16/5

Answer by BadLandsBarn(8)

Multiply by the reciprocal.
0.8 * 4/1 = 3.2

Answer by Alan3354(21583)

0.8/0.25 = 3.2/1 = 3.2
= 16/5